### v1

# Complex structures of the Gibbons-Hawking ansatz with infinite topological type

Abstract
In this paper, we study the complex structures of complete hyperkähler four-manifolds of infinite topological type arising from the Gibbons-Hawking ansatz. Under a natural genericity assumption on the monopole configuration, we show that the resulting manifold is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ defined by an explicit entire function, and we prove that this description persists for almost all complex structures in the associated $S^{2}$-family of compatible complex structures. When the genericity condition is relaxed, we further demonstrate that the manifold is biholomorphic to the minimal resolution of a singular surface. Thus, we partially extend LeBrun’s celebrated work [18] to the context of countably many monopoles.
Keywords. Gibbons-Hawking ansatz, infinite topological type, complex structures
1 Introduction
In their pioneering work, Gibbons and Hawking [9], [10] introduced a construction of four-dimensional hyperkähler metrics with a tri-Hamiltonian $S^{1}$-action, now known as the Gibbons-Hawking ansatz. Hitchin [14], through the twistor approach and the study of monopole moduli spaces, developed a coherent picture of the hyperkähler geometry behind these constructions. Kronheimer [17] described asymptotically locally Euclidean spaces arising as hyperkähler quotients. Bielawski [3] analyzed complete hyperkähler manifolds with local tri-Hamiltonian actions, placing a wide range of examples into a unified framework. In these works, the relevant configurations of monopole points are finite, and the resulting manifolds are of finite topological type. On the other hand, Anderson, Kronheimer and LeBrun [1] used Gibbons-Hawking metrics to construct complete Ricci-flat hyperkähler manifolds of infinite topological type. Goto [11] later gave a different viewpoint in terms of hyperkähler quotients, closely related to Kronheimer’s construction [17] of hyperkähler manifolds of type $A_{k}$. Subsequent work by Hattori [12] studied the volume growth and asymptotic geometry of these type $A_{\infty}$ manifolds. Closely related constructions appear in the theory of hypertoric manifolds, where Dancer and Swann analyzed some examples in [5, 6].
In this manuscript, inspired by the works of [1, 16, 18, 4], we investigate the complex structures of complete hyperkähler manifolds $M$ of infinite topological type obtained from the Gibbons-Hawking ansatz. Our first result (Theorem 1.1) gives an existence criterion for the Gibbons-Hawking potentials. The corresponding existence statement was already proved in work of Kalafat and Sawon [16] by an argument based on the maximum principle. We recover the criterion by a different method, using potential theory and the language of Riesz measures. Based on the method in LeBrun [18], we show that there exists a complex structure in the hyperkähler family on $M$ under which $M$ is biholomorphic to a hypersurface in $\mathbb{C}^{3}$ under a generic assumption (Theorem 1.3). LeBrun’s work treats the case where the circle action has finite fixed points and our description extends his picture to the case of a countable configuration of monopoles. After relaxing the genericity assumption, we further prove that the manifold can be realized as the minimal resolution of a singular hypersurface of a similar form (Theorem 1.4). In this way we relate the complex structures arising from the Gibbons–Hawking ansatz with infinite topological type directly to the geometry of the monopole set.
1.1 Background of Gibbons-Hawking ansatz
In this subsection we recall the Gibbons-Hawking construction and fix the particular hyperkähler manifold of infinite topological type that will be studied throughout the paper.
Let $U\subset\mathbb{R}^{3}$ be an open subset with coordinate $(x,y,z)$ and let $\pi_{U}:M_{U}\to U$ be a principal $S^{1}$-bundle. $V$ is a positive harmonic function defined on $U$ such that $\frac{1}{2\pi}\star dV$ represents the Chern class of this bundle, where $\star$ is the Hodge star operator with respect to the Euclidean metric $dg_{E}^{2}$ on $\mathbb{R}^{3}$. Then there is a connection $1$-form $\omega$ whose curvature is $\star dV$. The Gibbons-Hawking ansatz then produces a hyperkähler metric on $M_{U}$:
$g=\frac{1}{V}\omega^{2}+V\pi_{U}^{*}(dg_{E}^{2}).$
The corresponding Kähler form and complex structures are
$\Omega_{x}=dx\wedge\omega+Vdy\wedge dz,\ J_{x}:dx\mapsto V^{-1}\omega,\ dy\mapsto dz$
$\Omega_{y}=dy\wedge\omega+Vdz\wedge dx,\ J_{y}:dy\mapsto V^{-1}\omega,\ dz\mapsto dx$
$\Omega_{z}=dz\wedge\omega+Vdx\wedge dy,\ J_{z}:dz\mapsto V^{-1}\omega,\ dx\mapsto dy,$
and these satisfy the quaternionic relations
$J_{x}^{2}=J_{y}^{2}=J_{z}^{2}=-\text{Id},\ J_{x}J_{y}=-J_{y}J_{x}=J_{z}.$
Thus $g$ is a hyperkähler metric with a tri-Hamiltonian $S^{1}$-action generated by the circle fibres.
We now specialize to the class of Gibbons-Hawking metrics of infinite topological type introduced by Anderson-Kronheimer-LeBrun [1]. Consider a divergent sequence of distinct points $p_{j}=(x_{j},y_{j},z_{j})\in\mathbb{R}^{3},\ j\in\mathbb{N}^{+}$ and set $A=\{p_{j}:j\in\mathbb{N}^{+}\},\ U=\mathbb{R}^{3}\setminus A$. Let $\pi_{0}:M_{0}\to\mathbb{R}^{3}\setminus A$ be a principal $S^{1}$-bundle whose first Chern class has value 1-1 on a small sphere around each puncture $p_{j}$. More precisely, if $S_{j}^{2}$ is a sphere of sufficiently small radius enclosing only the point $p_{j}$, then the first Chern class $c_{1}$ of $\pi_{0}$ satisfies
$\int_{S^{2}_{j}}c_{1}=-1\text{ for every }j\in\mathbb{N}^{+}.$
[…]
$M:=M_{0}\cup\ \bigcup_{j=1}^{\infty}\hat{B_{j}}$$\sim$
[…]
Assume that the puncture points satisfy the summability condition
$\sum_{j=2}^{\infty}\frac{1}{{\left\|p_{1}-p_{j}\right\|}}<\infty.$ (1)
Then the function
$V(p):=\frac{1}{2}\sum_{j=1}^{\infty}\frac{1}{{\left\|p-p_{j}\right\|}}$ (2)
is a well-defined smooth harmonic function on $\mathbb{R}^{3}\setminus\{p_{j}\}$. One checks that the closed $2$-form $\frac{1}{2\pi}\star dV$ is a representative element of the Chern class $(-1,-1,\cdots)$ of the principal $S^{1}$-bundle. Hence there is a connection on $\pi_{0}$ with curvature $\star dV$. Let $\omega$ denote the connection form of this connection. The Gibbons-Hawking metric
$g=\frac{1}{V}\omega^{2}+V\pi_{0}^{*}(dg_{E}^{2})$
on $M_{0}$ extends smoothly across the added $S^{1}$-fixed points $\pi^{-1}(p_{j})\subset M$, and Anderson-Kronheimer-LeBrun [1] showed that the resulting metric $g$ on $M$ is complete and hyperkähler, while $M$ has infinite topological type.
In the sequel we shall mainly work with the complex structure $J:=J_{x}$. For later use it is convenient to describe it in slightly more concrete coordinates. Writing $\omega=dt+\theta$ for some $1$-form $\theta$ such that $d\theta=\star dV$, where $t$ is the angular coordinate along the $S^{1}$-fibres, the complex structure $J$ can be characterized by
$dx\mapsto V^{-1}(dt+\theta),$ (3)
$dy\mapsto dz.$ (4)
1.2 Main results
In this section, we present the main results derived from the hyperkähler manifolds of infinite topological type constructed in the previous section.
Throughout, let
  • $\{p_{j}\}_{j=1}^{\infty}\subset\mathbb{R}^{3}$ be a closed, discrete and countable set of points, with $p_{j}=(x_{j},y_{j},z_{j})$. We write $A=\{p_{j}:j\in\mathbb{N}^{+}\}$ and $U=\mathbb{R}^{3}\setminus A$.
  • $H:\mathbb{R}^{3}\to\mathbb{C}$ be the projection $H(x,y,z)=y+{\rm{i}}z$ and $a_{j}=H(p_{j})=y_{j}+{\rm{i}}z_{j}$.
We say that the configuration $\{p_{j}\}_{j=1}^{\infty}$ is generic if the projection
$H:\{p_{j}\}_{j=1}^{\infty}\to\mathbb{C}$
is injective, that is to say, $\{a_{j}\}_{j=1}^{\infty}$ are pairwise distinct.
Our first result is an existence criterion for the positive harmonic functions which arise as Gibbons-Hawking potential on $U$.
Theorem 1.1.
For $U=\mathbb{R}^{3}\setminus A$ where $A=\{p_{j}:j\in\mathbb{N}^{+}\}\subset\mathbb{R}^{3}$ is a closed, discrete and countable set, a nontrivial cohomology class
$\bm{e}=(e_{1},e_{2},\cdots)\in H^{2}(U,\mathbb{Z})$
can be represented in the form of $\star dV_{\bm{e}}]$ where $V_{\bm{e}}$ is some positive harmonic function on $U$, if and only if $e_{j}\leq 0$ for all $j\in\mathbb{N}^{+}$ and there exists a point $x\in U$ such that
$\sum_{j\geq 1}\frac{|e_{j}|}{|x-p_{j}|}<\infty.$
Remark 1.2.
In the concrete Gibbons-Hawking set-up in Section 1.1, the pricipal $S^{1}$-bundle $\pi_{0}:M_{0}\to\mathbb{R}^{3}\setminus\{p_{j}\}_{j=1}^{\infty}$ is chosen so that its first Chern class to be 1-1 on a small sphere enclosing each puncture $p_{j}$. Equivalently, the corresponding cohomology class $e=(e_{1},e_{2},\cdots)$ is fixed by the condition $e_{j}=-1$ for all $j\in\mathbb{N}^{+}$. Then we obtain a positive harmonic function $V_{\bm{e}}$ on $U=\mathbb{R}^{3}\setminus\{p_{j}\}_{j=1}^{\infty}$ of the form
$V_{\bm{e}}(x)=\sum_{j=1}^{\infty}\frac{1}{4\pi|x-p_{j}|}$
so that $\star dV_{\bm{e}}]=e$. This is exactly the potential appearing in (2) since $V=2\pi V_{\bm{e}}$. In particular, the requirement $e_{j}=-1$ for all $j$ ensures that the resulting hyperkähler metric extends smoothly across the added $S^{1}$-fixed points over $\{p_{j}\}_{j=1}^{\infty}$, producing the complete hyperkähler manifold $M$ of infinite topological type described in Section 1.1.
Our second main result describes $(M,J)$ under the generic hypothesis on the monopole points. For this we recall the standard Weierstrass primary factors
$E_{0}(u)=1-u,\ \ E_{j}(u)=(1-u)\exp\left(u+\frac{u^{2}}{2}+\cdots+\frac{u^{j}}{j}\right).$
and define the entire function
$P(u)=u^{\delta}\prod_{a_{j}\neq 0}E_{j}\left(u/a_{j}\right)$
where $\delta=0$ if $a_{j}\neq 0$ for all $j$ and $\delta=1$ if $0\in\{a_{j}\}$.
Theorem 1.3 (Generic case).
Suppose that the puncture points $\{p_{j}\}$ satisfy
$\sum_{j=2}^{\infty}\frac{1}{\|p_{1}-p_{j}\|}<\infty,$
and that the configuration is generic in the sense that the projections $a_{j}=H(p_{j})$ are pairwise distinct. If, in addition, the set $\{a_{j}\}_{j=1}^{\infty}$ has no accumulation point on $\mathbb{C}$, then $(M,J)$ is biholomorphic to the hypersurface in $\mathbb{C}^{3}$ given by
$u_{1}u_{2}=P(u_{3}).$
Moreover, we could prove in Section 3 that for almost all complex structures in the $S^{2}$-family, $M$ is likewise biholomorphic to a hypersurface of this specific form in $\mathbb{C}^{3}$.
When the generic assumption fails, several punctures may project to the same point on $\mathbb{C}$ under $H$. In this non-generic situation $M$ can no longer be described directly as a hypersurface in $\mathbb{C}^{3}$; instead, it is realized as the minimal resolution of a singular hypersurface.
Theorem 1.4 (Non-generic case).
For countable distinct points $p_{j}\in\mathbb{R}^{3},\ j\in\mathbb{N^{+}}$ and
$\sum_{j=2}^{\infty}\frac{1}{\|p_{1}-p_{j}\|}<\infty,$
set $a_{j}=H(p_{j})=y_{j}+{\rm{i}}z_{j}$ and let $\{b_{1},b_{2},\cdots\}$ be the set of distinct values of $\{a_{j}\}$. Assume that
  • (1).
    the set $\{b_{k}\}_{k=1}^{\infty}$ has no accumulation point on $\mathbb{C}$;
  • (2).
    for each $k\geq 1$, the following multiplicities are finite:
    $m_{0}:=\#\{j\in\mathbb{N}^{+}:a_{j}=0\}<\infty,$
    $m_{k}:=\#\{j\in\mathbb{N^{+}}:a_{j}=b_{k}\}<\infty.$
Relabel the points $\{p_{j}\}_{j}$ as $\{p_{k,l}\}_{k,l}$ so that $H^{-1}(b_{k})=\{p_{k,l}:1\leq l\leq m_{k}\}$ for all $k\in\mathbb{N}^{+}$. Using suitable primary factors, we define an entire function
$P(u):=u^{m_{0}}\prod_{k:b_{k}\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}}\right)\right)^{m_{k}}$
where the constants $c_{k}$ are properly chosen so that the infinite product converge uniformly on every compact subset of $\mathbb{C}$.
Let $S$ in $\mathbb{C}^{3}$ be the surface defined by
$u_{1}u_{2}=P(u_{3}).$
Then $(M,J)$ is biholomorphic to the minimal resolution of $S$.
1.3 Outline
The paper is structured as follows. In Section 2, we investigate the Riesz measure of the harmonic function $V$ and give the proof of Theorem 1.1. In Section 3, we discuss the complex structure of manifold $M$ under a generic assumption. By constructing explicit local charts and analyzing the equivalence relation, we prove that $(M,J)$ is biholomorphic to a hypersurface in $\mathbb{C}^{3}$. Furthermore, we show in Corollary 3.7 that this hypersurface equivalence holds for almost all complex structures in the $S^{2}$-family. Finally, in Section 4, we relax the generic assumption and discuss the non-generic case. Under some specified conditions we prove that $(M,J)$ is the minimal resolution of a hypersurface, as stated in Theorem 1.4.
2 Existence criteria for harmonic functions via Riesz measures
In this section we justify the choice of Gibbons-Hawking potential $V$ and give the proof of Theorem 1.1. The strategy is to apply the uniqueness of the Riesz measure associated to a subharmonic function and to compute this measure explicitly in terms of the monopole configuration.
Let $U=\mathbb{R}^{3}\setminus A$ where $A=\{p_{j}:j\in\mathbb{N}^{+}\}\subset\mathbb{R}^{3}$ is a closed, discrete and countable set. Then
$H^{2}(U,\mathbb{Z})=\prod_{j=1}^{\infty}\mathbb{Z}.$
We write a cohomology class ${\bm{e}}\in H^{2}(U,\mathbb{Z})$ as ${\bm{e}}=(e_{1},e_{2},\cdots)$.
When $e_{j}\leq 0\ \forall\ j\in\mathbb{N}^{+}$ and
$\sum_{j\geq 1}\frac{|e_{j}|}{|x-p_{j}|}<\infty,$
we can define
$V_{\bm{e}}(x)=\sum_{j=1}^{\infty}\frac{|e_{j}|}{4\pi|x-p_{j}|},\ x\in U.$
The series converges locally uniformly on $U$, and hence $V_{\bm{e}}$ is a well-defined harmonic function on $U$. To see this, set
$V_{\bm{e}}^{k}(x)=\sum_{j=1}^{k}\frac{|e_{j}|}{4\pi|x-p_{j}|}.$
Each $V_{\bm{e}}^{k}$ is a positive harmonic function on $U$, and the sequence $\{V_{\bm{e}}^{k}\}_{k\geq 1}$ is monotonically increasing. We invoke the following version of Harnack’s theorem (see [13, Theorem 1.20]).
Theorem 2.1 (Harnack’s Theorem).
For $\{u^{k}(x)\}$ a sequence of monotonically increasing harmonic functions on a connected open subset $U$ of $\mathbb{R}^{3}$, if there exists a point $q\in U$ such that the sequence $\{u^{k}(q)\}$ is bounded above, then there exists a harmonic function $u(x)$ on $U$ such that
$\lim_{k\to\infty}u^{k}(x)=u(x)$
uniformly on compact subsets of $U$.
In our case, choosing $x\in U$ with
$\sum_{j\geq 1}\frac{|e_{j}|}{|x-p_{j}|}<\infty,$
ensures that $\{V_{\bm{e}}^{k}(x)\}$ is uniformly bounded. By Harnack’s theorem, the sequence $\{V_{\bm{e}}^{k}\}_{k}$ converges locally uniformly on $U$ to a harmonic function, which we denote by $V_{\bm{e}}$. Thus $V_{\bm{e}}$ is positive and harmonic on $U$.
Moreover, for $j\in\mathbb{N}^{+}$,
$\int_{\partial B(p_{j},\epsilon_{j})}\star dV_{\bm{e}}=e_{j}$
for $\epsilon_{j}$ small enough, which implies that $\star dV_{\bm{e}}]={\bm{e}}$.
Conversely, we show that the conditions in Theorem 1.1 are necessary. Suppose $V_{\bm{e}}$ is a positive harmonic function such that
$[\star dV_{\bm{e}}]={\bm{e}}=(e_{1},e_{2},\cdots)\in H^{2}(U,\mathbb{Z}),$
[…]
If $V$ is a positive and harmonic function on $B(0,1)\setminus\{0\}\subset\mathbb{R}^{m}$ with $m>2$, then there exists a function $v$ harmonic on $B(0,1)$ and a constant $b\geq 0$ such that
$V(x)=v(x)+b|x|^{2-m}$ (5)
for all $x\in B(0,1)\setminus\{0\}$.
Therefore, if $V_{\bm{e}}$ is a positive harmonic function on $U$ such that $\star dV_{\bm{e}}]={\bm{e}}$, there exists a real harmonic function $V_{1}$ on $B(p_{1},\epsilon)$ and a real constant $b_{1}\geq 0$ such that
$V_{\bm{e}}=V_{1}+b_{1}\cdot\frac{1}{4\pi|x-p_{1}|},\ \ x\in B(p_{1},\epsilon)\setminus\{p_{1}\}$
for $\epsilon$ small enough. Furthermore,
$e_{1}=\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star dV_{\bm{e}}=\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star dV_{1}+\int_{\partial B(p_{1},{\frac{\epsilon}{2}})}\star d\left(b_{1}\cdot\frac{1}{4\pi|x-p_{1}|}\right)=-b_{1}.$
[…]
To extract the additional summability condition, we now use the language of Riesz measures. We recall two standard results from the theory of subharmonic functions [13, Theorem 3.9, Theorem 3.20].
Theorem 2.3.
If $u(x)$ is a subharmonic function on a connected open subset $U\subset\mathbb{R}^{m}$ with $m\geq 3$ and not identically -$\infty$, then there exists a unique Radon measure $\mu$ in $U$ satisfying that for any compact subset $E$ of $U$, there is a harmonic function $h(x)$ defined on the interior of $E$ and for $x$ where $h$ is defined,
$u(x)=\int_{E}K(x-\xi)d\mu+h(x)$ (6)
where
$K(x)=-|x|^{2-m}.$ (7)
The measure $\mu$ is the Riesz measure of $u$.
Theorem 2.4.
Given any Radon measure $\mu$ defined on $\mathbb{R}^{m}$ where $m\geq 3$, if $n(t)=\mu(\overline{B(0,t)})$, then $\mu$ is the Riesz measure of some subharmonic function $u(x)$ which is bounded from above on $\mathbb{R}^{m}$ if and only if
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}dt<\infty.$ (8)
Moreover, when this condition holds, the unique subharmonic function corresponding to $\mu$ is given by
$u(x)=C-\int_{\mathbb{R}^{m}}\frac{1}{|x-\xi|^{m-2}}d\mu.$ (9)
where $C$ is the least upper bound of $u(x)$.
Proof of Theorem 1.1.
Assume $V_{\bm{e}}$ is a positive harmonic function on $U$ with $\star dV_{\bm{e}}]={\bm{e}}$. We extend it to a subharmonic function on $\mathbb{R}^{3}$ by setting
$-V_{\bm{e}}(p_{i}):=\limsup_{x\to p_{i}}(-V_{\bm{e}})(x).$ (10)
In this way, the function $V_{\bm{e}}$ is upper semi-continuous on $\mathbb{R}^{3}$, satisfies $\infty\leq-V_{\bm{e}}(x)<\infty$, and obeys the mean-value inequality: for any point $x_{0}\in\mathbb{R}^{3}$ there exist a sufficiently small $r>0$ such that
$(-V_{\bm{e}})(x_{0})\leq\frac{1}{4\pi r^{2}}\int_{S(x_{0},r)}(-V_{\bm{e}})(x)d\sigma(x),$
where $d\sigma(x)$ denotes surface area on the sphere $S(x_{0},r)$. Define $u:=-V_{\bm{e}}$. Then $u$ is a subharmonic function on $\mathbb{R}^{3}$. Let $\mu$ be the Riesz measure of $u$ and set $n(t)=\mu(\overline{B(0,t)})$. We shall calculate $n(t)$.
[…]
$V_{\bm{e}}(x)=\frac{|e_{j}|}{4\pi|x-p_{j}|}+\widetilde{V_{j}}(x)$
where $\widetilde{V_{j}}$ is a harmonic function on $B(p_{j},\epsilon_{j})$. Thus $u(x)$ and $\frac{|e_{j}|}{4\pi|x-p_{j}|}$ determine the same Riesz measure on $B(p_{j},\epsilon_{j})$. In particular, for any $\varphi\in C_{0}^{\infty}(B(p_{j},\epsilon_{j}))$,
$\int_{B(p_{j},\epsilon_{j})}\varphi d\mu=-\int_{B(p_{j},\epsilon_{j})}\frac{|e_{j}|}{4\pi|x-p_{j}|}\cdot\Delta\varphi dx=|e_{j}|\cdot\varphi(p_{j}),$
so $\mu$ has an atom of mass $|e_{j}|$ at $p_{j}$.
Fix $t>0$ and set
$I_{t}=\{p_{j}:|p_{j}|\leq t\}=A\cap\overline{B(0,t)}.$
Then $I_{t}$ is a finite set. For $\epsilon>0$ sufficiently small, we can arrange that $I_{t}=A\cap B(0,t+\epsilon)$. Arguing as above, there exists harmonic function $v_{t}(x)$ on $B(0,t+\epsilon)$ such that
$V_{\bm{e}}(x)=\sum_{j:p_{j}\in I_{t}}\frac{|e_{j}|}{4\pi|x-p_{j}|}+v_{t}(x),\ \ x\in B(0,t+\epsilon)\setminus I_{t}$
and
$n(t)=\sum_{j:p_{j}\in I_{t}}|e_{j}|.$
Based on the argument above,
$n(t)=\mu(\overline{B(0,t)})=\sum_{j=1}^{\infty}|e_{j}|\cdot\chi_{\{|p_{j}|\leq t\}}(t),$
where
$\chi_{\{|p_{j}|\leq t\}}(t)=\left\{\begin{aligned} &1,&\text{ if }|p_{j}|\leq t,\\ &0,&\text{ if }|p_{j}|>t.\end{aligned}\right$.
To apply Theorem 2.4, we now evaluate the integral (8)
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}dt$
in the case $m=3$. Without loss of generality, we may reorder the points so that
$0<|p_{1}|\leq|p_{2}|\leq|p_{3}|\leq\cdots$
which preserves the set $A$. Assume that there exists some index $i$ such that $|p_{i}|\leq 1$. Let $N$ be the index satisfying
$|p_{1}|\leq\cdots\leq|p_{N}|\leq 1<|p_{N+1}|\leq\cdots.$
Then
$\int_{1}^{\infty}\frac{n(t)}{t^{2}}dt$ $=\int_{1}^{|p_{N+1}|}\frac{1}{t^{2}}\left(\sum_{i=1}^{N}|e_{i}|\right)dt+\sum_{i\geq N+1}^{\infty}\int_{|p_{i}|}^{|p_{i+1}|}\frac{1}{t^{2}}\left(\sum_{j=1}^{i}|e_{j}|\right)dt$
$=\left(1-\frac{1}{|p_{N+1}|}\right)\left(\sum_{i=1}^{N}|e_{i}|\right)+\sum_{i\geq N+1}\left(\frac{1}{|p_{i}|}-\frac{1}{|p_{i+1}|}\right)\left(\sum_{j=1}^{i}|e_{j}|\right)$
$=\sum_{i=1}^{N}|e_{i}|+\sum_{i\geq N+1}\frac{|e_{i}|}{|p_{i}|}.$
If instead $|p_{j}|>1$ for all $j\in\mathbb{N}^{+}$, then the same computation yields
$\int_{1}^{\infty}\frac{n(t)}{t^{m-1}}dt$ $=\sum_{j=1}^{\infty}\frac{|e_{j}|}{|p_{j}|}.$
Finally, by a simple application of the triangle inequality, the convergence of $\sum_{j}\frac{|e_{j}|}{|p_{j}|}$ is equivalent to the existence of a point $x\in U$ such that
$\sum_{j=1}^{\infty}\frac{|e_{j}|}{|x-p_{j}|}<\infty.$
Thus Theorem 2.4 shows that the summability condition in Theorem 1.1 is also necessary, completing the proof.
3 The complex structure of $M$ under generic assumptions
This section is devoted to describing the complex structure $J$ defined in (3) on the hyperkähler manifold $M$ of infinite topological type constructed through the Gibbons-Hawking ansatz in Section 1.1. Under the generic assumption we prove that $(M,J)$ is biholomorphic to an explicit hypersurface in $\mathbb{C}^{3}$. Our approach builds on the work of LeBrun[18], but is adapted to the case where the $S^{1}$-action has countably many fixed points.
3.1 A holomorphic $\mathbb{C}^{*}$ action on $M$
Let $\frac{\partial}{\partial t}$ denote the vector field generating the $S^{1}$-action along the fibres of $\pi:M\to\mathbb{R}^{3}$. Since the hyperkähler structure is $S^{1}$-invariant, we have $\mathcal{L}_{\frac{\partial}{\partial t}}J=0$, so $\frac{\partial}{\partial t}$ is a real holomorphic vector field with respect to $J$. Its $(1,0)$-part is
$\left(\frac{\partial}{\partial t}\right)^{1,0}=\frac{1}{2}\left(\frac{\partial}{\partial t}-{\rm{i}}J\frac{\partial}{\partial t}\right).$
We now define a holomorphic vector field
$\xi:=-{\rm{i}}\left(\frac{\partial}{\partial t}\right)^{1,0}=-\frac{{\rm{i}}}{2}\left(\frac{\partial}{\partial t}-{\rm{i}}J\frac{\partial}{\partial t}\right)=\frac{1}{2}\left(V^{-1}\hat{\frac{\partial}{\partial x}}-{\rm{i}}\frac{\partial}{\partial t}\right)$
where $\hat{\frac{\partial}{\partial x}}$ is the horizontal lift of $\frac{\partial}{\partial x}$ with respect to the connection $\omega$. The last equality follows from the explicit description of $J$ in (3).
Note that $V^{-1}\hat{\frac{\partial}{\partial x}}$ is the horizontal lift of the vector field $V^{-1}\frac{\partial}{\partial x}$ on $\mathbb{R}^{3}$. Hence the completeness of flow generated by $V^{-1}\hat{\frac{\partial}{\partial x}}$ is equivalent to that of flow generated by $V^{-1}\frac{\partial}{\partial x}$ on $\mathbb{R}^{3}$. Consider an integral curve $(x(s),y,z)$ of $V^{-1}\frac{\partial}{\partial x}$. Along such a curve we have
$\frac{dx(s)}{ds}=V^{-1}(x,y,z),$
and we have
$\int_{x(s_{1})}^{x(s_{2})}V(x,y,z)dx=\int_{s_{1}}^{s_{2}}ds=s_{2}-s_{1}.$
Since $V$ is positive and smooth on $U$, these integral curves extend for all $s\in\mathbb{R}$, and the flow of $V^{-1}\frac{\partial}{\partial x}$ on $\mathbb{R}^{3}$ is complete. It follows that the real and imaginary parts of $\xi$ are complete vector fields, and hence $\xi$ generates a holomorphic $\mathbb{C}$-action with complete complex flow. We denote this flow by
$\Phi_{\lambda}:M\to M,\ \lambda\in\mathbb{C}.$
Writing $\lambda=t+{\rm{i}}s$, the flow may be decomposed as
$\Phi_{t+is}(p)=\phi_{t}\circ\psi_{s}(p)=\psi_{s}\circ\phi_{t}(p)$
where the real flows $\phi_{t}$ and $\psi_{s}$ are defined by
$\frac{\partial\phi_{t}}{\partial t}=V^{-1}\hat{\frac{\partial}{\partial x}}=-J\frac{\partial}{\partial t},$
$\frac{\partial\psi_{s}}{\partial s}=JV^{-1}\hat{\frac{\partial}{\partial x}}=\frac{\partial}{\partial t}.$
In particular, $\psi_{s}$ coincides with the original $S^{1}$-action and $\psi_{s}=\psi_{s+2\pi}$. Thus $\Phi_{t+{\rm{i}}s}=\Phi_{t+{\rm{i}}s+2\pi{\rm{i}}}$. The $\mathbb{C}$-action on $M$ descends via the exponential map to a holomorphic $\mathbb{C}^{*}$-action on $M$. Concretely,
$\mathbb{C}^{*}\times M\to M,\ (w,p)\mapsto\Phi_{\log w}(p).$
Next we describe the orbits of this $\mathbb{C}^{*}$-action. The holomorphic function
$H:M$ $\to\mathbb{C},$
$(x,y,z,t)$ $\mapsto y+{\rm{i}}z$
is the natural extension of the projection $H:\mathbb{R}^{3}\to\mathbb{C}$ introduced in Section 1.2. For $u\in\mathbb{C}$, let
$L_{u}:=\{(x,y,z)\in\mathbb{R}^{3}:y+{\rm{i}}z=u\}$
denote a straight line in $\mathbb{R}^{3}$ parallel to $x$-axis and set $C_{u}:=\pi^{-1}(L_{u})\subset M$. Then we have:
  • If $u\neq y_{j}+{\rm{i}}z_{j}$, $C_{u}$ is a single orbit.
  • If $u=y_{j}+{\rm{i}}z_{j}$ for some $j$, $C_{u}$ is the union of three orbits: the fixed point $\{q_{j}:=\pi^{-1}(p_{j})\}$, together with two open orbits $\pi^{-1}(\{(x,y_{j},z_{j}):x>x_{j}\})$ and $\pi^{-1}(\{(x,y_{j},z_{j}):x<x_{j}\})$.
Consider
$\Psi:=-{\rm{i}}(dy\wedge\omega+Vdz\wedge dx)+(dz\wedge\omega+Vdx\wedge dy)$
which is a parallel $(2,0)$-form preserved by $\xi$. The holomorphic function $H:M\to\mathbb{C},\ q=(x,y,z,t)\mapsto y+{\rm{i}}z$ satisfies
$dH=\iota_{\xi}\Psi.$
$M_{0}=M\setminus\{\pi^{-1}(\{p_{j}\})\}$ is the open subset of $M$ where the $S^{1}$-action is free. We can cover $M_{0}$ by two open subsets:
$M^{+}:=\pi^{-1}\left(\{u\neq y_{j}+{\rm{i}}z_{j}\text{ for all }j\}\cup\bigcup_{j}\{u=y_{j}+{\rm{i}}z_{j}\text{ and}\ x>x_{j}\}\right)$
$M^{-}:=\pi^{-1}\left(\{u\neq y_{j}+{\rm{i}}z_{j}\text{ for all }j\}\cup\bigcup_{j}\{u=y_{j}+{\rm{i}}z_{j}\text{ and}\ x<x_{j}\}\right).$
Then $H:M^{\pm}\to\mathbb{C}$ are holomorphic principal $\mathbb{C}^{*}$-bundles. $\mathbb{C}$ is Stein manifold. Since $\mathbb{C}^{*}$ deformation-retracts to $S^{1}$, we have
$[\mathbb{C},K(\mathbb{Z},2)]\cong H^{2}(\mathbb{C},Z)=0.$
Hence the principal $\mathbb{C}^{*}$-bundles over $\mathbb{C}$ are all topologically trivial. Then by the Oka-Grauert Principle, [8, Theorem 8.2.1], $M^{\pm}$ are both biholomorphic to $\mathbb{C}\times\mathbb{C}^{*}$ as holomorphic principal $\mathbb{C}^{*}$-bundles over $\mathbb{C}$.
Theorem 3.1 (Oka-Grauert Principle).
For Stein manifold $X$ and every complex Lie group $G$, the holomorphic and the topological classes of principal $G$-bundles over $X$ are in one-to-one correspondence.
Hence there exist biholomorphisms
$H^{\pm}:M^{\pm}$ $\to\mathbb{C}\times\mathbb{C}^{*}$
$q$ $\mapsto(u,v)_{\pm}$
Under these identifications, the vector field $\xi$ is mapped to $v\frac{\partial}{\partial v}$ and the $2$-form $\Psi$ is mapped to $\frac{dv\wedge du}{v}.$ The induced $\mathbb{C}^{*}$-action is therefore
$\lambda\cdot(u,v)=(u,\lambda v),\ \lambda\in\mathbb{C}^{*}.$
3.2 The equivalence relation
Now $M_{0}$ is biholomorphic to
[…]
where $\sim$ denotes the equivalence
$(u,v)_{+}\sim(u,f(u)v)_{-}$
for some holomorphic function $f:\mathbb{C}\setminus\{y_{j}+{\rm{i}}z_{j}\}\to\mathbb{C}^{*}$.
Lemma 3.2.
For holomorphic $g:\mathbb{C}\to\mathbb{C}^{*}$, the two gluings defined by $f$ and $fg$ give biholomorphic complex manifolds.
Proof.
Define holomorphic changes of trivialization
$\psi^{+}:(u,v)_{+}$ $\mapsto(u,v)_{+},$
$\psi^{-}:(u,v)_{-}$ $\mapsto(u,g(u)v)_{-}.$
Then the equivalence relation would be
$\psi^{-}\circ H^{-}\circ(H^{+})^{-1}\circ(\psi^{+})^{-1}=(u,f(u)g(u)v).$
Therefore, the complex structure can be decided only by behaviors of $f$ near the punctures. ∎
Lemma 3.3.
On the overlap $M^{+}\cap M^{-}$, the equivalence relation can be given by
$(u,v)_{+}\sim\left(u,\frac{v}{P(u)}\right)_{-}$
where
$P(u)=u^{\delta}\prod_{a_{j}\neq 0}E_{j}\left(u/a_{j}\right).$
$\delta=0$ if $a_{j}\neq 0$ for all $j$ and $\delta=1$ if $0\in\{a_{j}\}$.
We shall apply the following complex Morse Lemma, [7, Proposition 3.15]
Lemma 3.4.
Let $M$ be a $2$-dimensional complex manifold, $H:M\to\mathbb{C}$ a holomorphic function and $q$ a nondegenerate critical point of $H$. Then there is a local coordinate system $(w_{1},w_{2})$ in a neighbourhood $V$ of $q$ such that on $V$ we have
$H(w_{1},w_{2})=H(q)+w_{1}w_{2},\ \ w_{1}(q)=w_{2}(q)=0.$
Proof of Lemma 3.3.
Now $H:M\to\mathbb{C}$ has nondegenerate critical point at the isolated points $q_{j}=\pi^{-1}(p_{j})$ for $p_{j}=(x_{j},y_{j},z_{j})$. Indeed, we have
$dH_{p_{j}}=\iota_{\xi}\Psi|_{p_{j}}=0$
and the Hessian of $H$ at $q_{j}$ can be expressed as
$\nabla dH(v,w)=\left(\nabla\iota_{\xi}\Psi(v,w)\right)=\left(\nabla_{v}\Psi(\xi,w)\right)-\Psi(\xi,\nabla_{v}w)=\Psi(\nabla_{v}\xi,w)$
Note that the $\mathbb{C}^{*}$-action generated by $\xi$ has isolated fixed points $\{p_{j}\}$, the linear map $\nabla\xi|_{q_{j}}:T_{q_{j}}M\to T_{q_{j}}M$ is invertible. As $\Psi$ is nondegenerate, it follows that the Hessian of $H$ is non-degenerate at each $q_{j}$. By the Morse Lemma each $q_{j}$ has a neighbourhood $U_{j}$ with a holomorphic chart on which we could find complex coordinates $(w^{j}_{1},w^{j}_{2})$ such that
$\varphi^{j}=(w^{j}_{1},w^{j}_{2}):U_{j}\to V_{j}\subset\mathbb{C}^{2},\ \varphi^{j}(q_{j})=0,$
$H\circ(\varphi^{j})^{-1}(w^{j}_{1},w^{j}_{2})=y_{j}+{\rm{i}}z_{j}+w^{j}_{1}w^{j}_{2}=a_{j}+w^{j}_{1}w^{j}_{2}.$
Where no confusion can arise, we drop the index $j$ for brevity. On $(U_{j},\varphi^{j})$, we have
$\Psi=\psi(w_{1},w_{2})dw_{1}\wedge dw_{2}$
for some nonvanishing holomorphic function $\psi(w_{1},w_{2})$. Then $\xi=a\frac{\partial}{\partial w_{1}}+b\frac{\partial}{\partial w_{2}}$ can be determined by
$\iota_{\xi}\Psi=dH=w_{1}dw_{2}+w_{2}dw_{1}=\iota_{\xi}(\psi dw_{1}\wedge dw_{2}),$
which implies that
$\xi=\psi^{-1}(w_{1},w_{2})\left(w_{1}\frac{\partial}{\partial w_{1}}-w_{2}\frac{\partial}{\partial w_{2}}\right).$
Since $\xi$ generates $\mathbb{C}^{*}$-action,
$\int_{0}^{2\pi}\psi(e^{{\rm{i}}\theta}w_{1},e^{-{\rm{i}}\theta}w_{2})d\theta=2\pi.$
The complex flow of $\xi$ with respect to $\lambda=s+it$
$\frac{dw_{1}}{d\lambda}=\psi^{-1}w_{1}$
$\frac{dw_{2}}{d\lambda}=-\psi^{-1}w_{2}$
Integrate $\lambda$ along a closed circle around $0\in\mathbb{C}$,
$\oint d\lambda=2\pi{\rm{i}}=\oint\psi\frac{dw_{1}}{w_{1}}={\rm{i}}\int_{0}^{2\pi}\psi(e^{{\rm{i}}\theta}w_{1},e^{-{\rm{i}}\theta}w_{2})d\theta.$
Consider the expansion of $\psi$
$\psi=\sum_{k,l=0}^{\infty}a_{kl}\cdot w_{1}^{k}w_{2}^{l},\ a_{00}=1,\ a_{jj}=0,\ \ \forall\ j>0.$
One can solve the equation
$\left(w_{1}\frac{\partial}{\partial w_{1}}-w_{2}\frac{\partial}{\partial w_{2}}\right)\alpha=\psi-1$
by
$\alpha=\sum_{k\neq l}\frac{a_{kl}}{k-l}w_{1}^{k}w_{2}^{l}.$
Now set
$s^{j}_{1}:=e^{\alpha}w_{1}^{j},\ s^{j}_{2}:=e^{-\alpha}w_{2}^{j}.$
Then $(s^{j}_{1},s^{j}_{2})$ gives a chart around $q_{j}$ satisfying
$H=s_{1}^{j}s_{2}^{j}+a_{j},\ a_{j}=y_{j}+{\rm{i}}z_{j}.$
and we have
$\xi=s_{1}\frac{\partial}{\partial s_{1}}-s_{2}\frac{\partial}{\partial s_{2}}.$
The flow generated by $\xi$ is given by
$\Phi_{\lambda}(s_{1},s_{2})=(e^{\lambda}s_{1},e^{-\lambda}s_{2}),\ \lambda\in\mathbb{C}.$
Claim: Near the critical point $q_{j}$ where $u=H\neq y_{j}+{\rm{i}}z_{j}$, the equivalence relation is
$f_{j}(u)=\frac{\epsilon^{2}}{u-(y_{j}+{\rm{i}}z_{j})}=\frac{\epsilon^{2}}{u-a_{j}}$
where $\epsilon$ is a non-vanishing complex constant. Under coordinate transformation, the equivalence function can always be represented by $f_{j}(u)g_{j}(u)$ where $g_{j}$ is a holomorphic non-vanishing function on $U_{j}$.
Indeed, we have
$U_{j}\cap M^{+}=\{s_{1}^{j}s_{2}^{j}\neq 0\}\sqcup(\{s_{1}^{j}=0,s_{2}^{j}\neq 0\}\cap M^{+})\sqcup(\{s_{1}^{j}\neq 0,s_{2}^{j}=0\}\cap M^{+}),$
$U_{j}\cap M^{-}=\{s_{1}^{j}s_{2}^{j}\neq 0\}\sqcup(\{s_{1}^{j}=0,s_{2}^{j}\neq 0\}\cap M^{-})\sqcup(\{s_{1}^{j}\neq 0,s_{2}^{j}=0\}\cap M^{-}).$
Since $U_{j}\cap M^{+}\setminus\{s_{1}^{j}s_{2}^{j}\neq 0\}$ is $U_{j}\cap M^{+}\setminus\{u\neq a_{j}\}=U_{j}\cap\{u=a_{j},\ x>x_{j}\}$ which contains only one orbit, it must be one of the two cases, either $\{s_{1}^{j}=0,s_{2}^{j}\neq 0\}\cap M^{+}$ or $\{s_{1}^{j}\neq 0,s_{2}^{j}=0\}\cap M^{+}$.
Without loss of generality, we may assume the former one. $(U_{j}\cap M^{+})\cup(U_{j}\cap M^{-})$ has been mapped onto $V_{j}\setminus\{0\}\subset\mathbb{C}^{2}$, then we can exactly pick standard section $\sigma_{+}$ for $U_{j}\cap M^{+}$ and $\sigma_{-}$ for $U_{j}\cap M^{-}$:
$\sigma_{+}=(s_{1}(u),s_{2}(u))=(\epsilon,\frac{u-a_{j}}{\epsilon}),$
$\sigma_{-}=(s_{1}(u),s_{2}(u))=(\frac{u-a_{j}}{\epsilon},\epsilon),$
for some fixed non-zero constant $\epsilon\in\mathbb{C}$. The local trivializations for $H:M^{\pm}\to\mathbb{C}$ restricted on $U_{j}$ can be written as
$h_{j}^{+}:V_{j}^{+}$ $\to M^{+}\cap U_{j}$
$(u,v^{+})$ $\mapsto(s_{1},s_{2})=\Phi_{\log v^{+}}(\epsilon,\frac{u-a_{j}}{\epsilon})=(\epsilon v^{+},\frac{u-a_{j}}{\epsilon v^{+}}).$
$h_{j}^{-}:V_{j}^{-}$ $\to M^{-}\cap U_{j}$
$(u,v^{-})$ $\mapsto(s_{1},s_{2})=\Phi_{\log v^{-}}(\frac{u-a_{j}}{\epsilon},\epsilon)=(\frac{u-a_{j}}{\epsilon}v^{-},\frac{\epsilon}{v^{-}}).$
where $V_{j}^{+}\subset(H(U_{j})-\{a_{j}\})\times\mathbb{C}^{*}$, $V_{j}^{-}\subset(H(U_{j})-\{a_{j}\})\times\mathbb{C}^{*}$ are two suitable open subsets.
And the equivalence relation is decided by
$(h_{j}^{-})^{-1}\circ h_{j}^{+}(u,v^{+})=(u,\frac{\epsilon^{2}}{u-a_{j}}v^{+}).$
We now relate this local description to the Weierstrass product introduced in Section 1.2. Recall that under the generic assumption we have a countable set $\{a_{j}\}$ of pairwise distinct points of $\mathbb{C}$ without accumulation, and we define
$P(u)=u^{\delta}\prod_{a_{j}\neq 0}E_{j}\left(u/a_{j}\right)$
where $\delta=0$ if $a_{j}\neq 0$ for all $j$ and $\delta=1$ if $0\in\{a_{j}\}$, and where $E_{j}$ are the standard Weierstrass primary factors. Then $P(u)$ is an entire function with simple zeros exactly at the points $a_{j}$.
If the equivalence relation of $M_{0}$ is $f$, then $f$ must have exactly simple poles at $a_{j}$ and no other singularities. Now $f$ and $\frac{1}{P}$ have the same zeros and poles of same multiplicities at $\{a_{j}\}$, then the quotient
$g(u):=f(u)P(u)$
extends to an entire holomorphic function $g:\mathbb{C}\to\mathbb{C}^{*}$. From Lemma 3.2 and based on the discussions above, we have coordinates $(u,v_{\pm})$ on $M^{\pm}$ and the equivalence relation is given by $\frac{1}{P(u)}$.
We now give the proof of our second main result.
Proof of Theorem 1.3.
Consider the hypersurface
$\Sigma:=\{(u_{1},u_{2},u_{3})\in\mathbb{C}^{3}:u_{1}u_{2}=P(u_{3})\}.$
Suppose $\sigma:=\{(0,0,y_{j}+{\rm{i}}z_{j}):j\in\mathbb{N}^{+}\}\subset\mathbb{C}^{3}$ and define
$\chi:M_{0}$ $\to\Sigma\setminus\sigma,$
$(u,v_{+})$ $\mapsto(u_{1},u_{2},u_{3})=(\frac{P(u)}{v_{+}},v_{+},u),$
$(u,v_{-})$ $\mapsto(u_{1},u_{2},u_{3})=(\frac{1}{v_{-}},P(u)v_{-},u).$
Then $\chi(M^{+})=\Sigma^{+}=\{u_{2}\neq 0\}\cap\Sigma\setminus\sigma$, $\chi(M^{-})=\Sigma^{-}=\{u_{1}\neq 0\}\cap\Sigma\setminus\sigma$. $\chi$ is well-defined and holomorphic on $M_{0}$.
Conversely, on $\Sigma\setminus\sigma$ we define
$(u_{1},u_{2},u_{3})\mapsto\left\{\begin{aligned} (u_{3},u_{2})_{+}\ &\text{ if }u_{2}\neq 0,\\ \left(u_{3},\frac{1}{u_{1}}\right)_{-}\ &\text{ if }u_{1}\neq 0,\end{aligned}\right$.
which gives a holomorphic inverse to $\chi$. Hence $\chi$ is a biholomorphism from $M_{0}$ to $\Sigma\setminus\sigma$ and locally maps $U_{j}\setminus\{q_{j}\}$ to the intersection of $\Sigma\setminus\sigma$ and a neighbourhood of $(0,0,y_{j}+{\rm{i}}z_{j})$ in $\Sigma$.
By the following Hartogs’s theorem [15, Theorem 2.3.2], $\chi$ extends uniquely to a biholomorphism from $M$ to $\Sigma$. ∎
Theorem 3.5 (Hartogs’s extension theorem).
For an open set $U\subset\mathbb{C}^{n},\ n\geq 2$, any holomorphic function on $U\setminus\{q\}$ which is a set punctured at one point $q\in U$ can be extended to a holomorphic function on $U$.
Remark 3.6.
Theorem 1.3 recovers, as a special case, the example of Anderson-Kronheimer-LeBrun [1] in which $\{p_{j}\}_{j}=\{(0,0,z_{j})\in\mathbb{R}^{3}\}_{j}$ with $\sum_{j=1}^{\infty}\frac{1}{|z_{j}|}<\infty$.
In this situation one has $a_{j}={\rm{i}}z_{j}$ and the Weierstrass product reduces to
$P(u)=\prod_{j=1}^{\infty}\left(1-\frac{u_{3}}{{\rm{i}}z_{j}}\right).$
So $(M,J)$ is biholomorphic to
$u_{1}u_{2}=\prod_{j=1}^{\infty}\left(1-\frac{u_{3}}{z_{j}}\right).$
So far we have worked with the complex structure $J=J_{x}:dx\mapsto V^{-1}(dt+\theta),\ dy\mapsto dz$, for which the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to the $x$-axis. Recall we have
$J_{y}:dy\mapsto V^{-1}(dt+\theta),\ dz\mapsto dx,$
$J_{z}:dz\mapsto V^{-1}(dt+\theta),\ dx\mapsto dy.$
Together, these three complex structures satisfy the quaternionic identities
$J_{x}^{2}=J_{y}^{2}=J_{z}^{2}=-\text{Id},\ J_{x}J_{y}=-J_{y}J_{x}=J_{z}.$
Then for each direction $v=(v_{1},v_{2},v_{3})\in S^{2}\subset\mathbb{R}^{3}$, we obtain a compatible complex structure
$J_{v}:=v_{1}J_{x}+v_{2}J_{y}+v_{3}J_{z}$
for which the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to $v$. The projection to the plane $v^{\perp}$
$\Pi_{v}=\text{Id}-vv^{T}:\mathbb{R}^{3}\to v^{\perp}\cong\mathbb{C}.$
plays the role of $H$ in this setting.
We now show that for almost such directions $v$ the previous construction applies.
Corollary 3.7.
For $v\in S^{2}$, define
$a_{j}(v):=\Pi_{v}(p_{j})\text{ and }P_{v}(u):=u^{\delta(v)}\prod_{a_{j}(v)\neq 0}E_{j}\left(\frac{u}{a_{j}(v)}\right).$
where $\delta(v)=0$ if $a_{j}(v)\neq 0$ for all $j$ and $\delta(v)=1$ if otherwise. Then for almost every $v\in S^{2}$, the points $\{a_{j}(v)\}$ are pairwise distinct, have no accumulation point in $\mathbb{C}$, and the conclusion of Theorem 1.3 holds with $H$ replaced by $\Pi_{v}$.
In particular, for such $v$, $(M,J_{v})$ is biholomorphic to the hypersurface
$u_{1}u_{2}=P_{v}(u_{3})\subset\mathbb{C}^{3}.$
Proof.
[…]
For any fixed $n$ we have
[…]
Since
[…]
we have
[…]
By the Borel-Cantelli lemma, for $A_{n}:=\{v\in S^{2}:\ v\in A_{j,n}\text{ for infinitely many }j\}=\limsup_{j\to\infty}A_{j,n}$,
$m(A_{n})=0.$
Then the probability that $\Pi_{v}(p_{j})$ has some accumulation point is
[…]
$=$ $0.$
Secondly, the set
[…]
is a countable set. Hence for every $v\in S^{2}\setminus B$, $\Pi_{v}$ is injective on $\{p_{j}\}$. Then Theorem 1.3 completes the proof.
4 Non-injective projection and the minimal resolution
While the previous discussion focused on the case where $H$ is bijective from $\{p_{j}\}_{j}$ to their images, where the complex structure of the infinite topological type hyperkähler manifold $M$ is realized by a hypersurface in $\mathbb{C}^{3}$, the aim of this section is to relax this injectivity assumption and to analyse the resulting non-generic configurations. When $H$ fails to be one-to-one, $M$ should then be interpreted not as the hypersurface itself, but as its minimal resolution of singularities.
Before turning to the general countable configuration, we first recall a simple model case in which exactly two fixed points lie in the same image of $H$. This example, due to Chen-Chen [4], already illustrates how surface singularities of type $A_{1}$ arise in this setting and how they are resolved. It will serve as a concrete motivation for the general statement of Theorem 1.4.
Example 4.1 (Chen-Chen [4]).
Without loss of generality, assume that there are only two fixed points under the $S^{1}$ action: $q_{1}=\pi^{-1}(p_{1}),\ q_{2}=\pi^{-1}(p_{2})$ where $p_{1}=(0,0,0)$ and $p_{2}=(1,0,0)$. In this situation, Chen-Chen [4] has shown that the resulting hyperkähler manifold $M$ is biholomorphic to the minimal resolution of the hypersurface of $\mathbb{C}^{3}$ defined by $u_{1}u_{2}=u_{3}^{2}$.
Define
$M_{1}$ $:=\pi^{-1}\left(\{u\neq 0\}\cup\{u=0\text{ and }x<0\}\right),$
$M_{2}$ $:=\pi^{-1}\left(\{u\neq 0\}\cup\{u=0\text{ and }0<x<1\}\right),$
$M_{3}$ $:=\pi^{-1}\left(\{u\neq 0\}\cup\{u=0\text{ and }x>1\}\right).$
Then $M_{1}\cup M_{2}\cup M_{3}$ is an open cover of $M_{0}:=M\setminus\{\pi^{-1}(\{p_{1},p_{2}\})\}$ and each $M_{j},\ j\in\{1,2,3\}$ is a holomorphic principal $\mathbb{C}^{*}$-bundle under $H:M_{j}\to\mathbb{C}$. $M_{0}$ is biholomorphic to
$(\mathbb{C}\times\mathbb{C}^{*})_{M_{1}}\amalg(\mathbb{C}\times\mathbb{C}^{*})_{M_{2}}\amalg(\mathbb{C}\times\mathbb{C}^{*})_{M_{3}}/\sim$
where the equivalence relation $\sim$ is generated by
$(u,v)_{M_{1}}\sim(u,\frac{v}{u})_{M_{2}}\sim(u,\frac{v}{u^{2}})_{M_{3}}.$
Let $S$ be the hypersurface in $\mathbb{C}^{3}$ defined by the following equation
$u_{1}u_{2}=u_{3}^{2},$
and let $Bl_{0}(S)$ denoted the blowing-up of $S$ at point $(0,0,0)\in\mathbb{C}^{3}$:
$Bl_{0}(S)$ $=\{((u_{1},u_{2},u_{3}),[U_{1},U_{2},U_{3}]):u_{1}U_{2}=u_{2}U_{1},\ u_{1}U_{3}=u_{3}U_{1},\ u_{2}U_{3}=u_{3}U_{2},\ U_{1}U_{2}=U_{3}^{2}\}$
$\subset S\times\mathbb{P}^{2}.$
Remove the two points of the exceptional divisor over the origin by setting
$Bl_{0}(S)^{\prime}:=Bl_{0}(S)\setminus\{\left((0,0,0),[1:0:0]\right),\left((0,0,0),[0:1:0]\right)\}.$
On each $M_{j}$ we choose local coordinates $(u,v_{j})$ and define a map $\chi_{0}:M_{0}\to Bl_{0}(S)^{\prime}$ by
$\chi_{0}|_{M_{1}}(u,v_{1})$ $=\left(\left(\frac{u^{2}}{v_{1}},v_{1},u\right),\left[\left(\frac{u}{v_{1}}\right)^{2}:1:\frac{u}{v_{1}}\right]\right),$
$\chi_{0}|_{M_{2}}(u,v_{2})$ $=\left(\left(\frac{u}{v_{2}},uv_{2},u\right),\left[\frac{1}{v_{2}}:v_{2}:1\right]\right),$
$\chi_{0}|_{M_{3}}(u,v_{3})$ $=\left(\left(\frac{1}{v_{3}},u^{2}v_{3},u\right),\left[1:\left(uv_{3}\right)^{2}:uv_{3}\right]\right).$
These expressions are compatible on overlaps, so $\chi_{0}$ is well-defined and holomorphic.
Conversely, consider the open subsets
$\Sigma_{1}:=\{u_{2}\neq 0\}\cap Bl_{0}(S)^{\prime},$
$\Sigma_{2}:=\{U_{3}\neq 0\}\cap Bl_{0}(S)^{\prime},$
$\Sigma_{3}:=\{u_{1}\neq 0\}\cap Bl_{0}(S)^{\prime}.$
Define
$\Theta_{0}|_{\Sigma_{1}}\left(\left(u_{1},u_{2},u_{3}\right),\left[U_{1}:U_{2}:U_{3}\right]\right)=(u_{3},u_{2})_{M_{1}},$
$\Theta_{0}|_{\Sigma_{2}}\left(\left(u_{1},u_{2},u_{3}\right),\left[U_{1}:U_{2}:U_{3}\right]\right)=(u_{3},\frac{U_{2}}{U_{3}})_{M_{2}},$
$\Theta_{0}|_{\Sigma_{3}}\left(\left(u_{1},u_{2},u_{3}\right),\left[U_{1}:U_{2}:U_{3}\right]\right)=(u_{3},\frac{1}{u_{1}})_{M_{3}}.$
Again, one checks that $\Theta_{0}$ is well-defined and holomorphic, and that $\Theta_{0}$ and $\chi_{0}$ are inverse to each other. Hence we have the biholomorphism between $M_{0}$ and $Bl_{0}(S)^{\prime}$. By the Hartogs’s theorem, this biholomorphism extends across the two missing points to a biholomorphism $M\cong Bl_{0}(S)$. So in the two-centre case $(M,J)$ is exactly the minimal resolution of the singular surface $S$.
We now generalize this picture to the case of countably many monopole points, where several monopoles may project to the same point under $H$.
Let $\{p_{j}\}_{j\in\mathbb{N}^{+}}\subset\mathbb{R}^{3}$ be a countable set of distinct points such that
$\sum_{j=2}^{\infty}\frac{1}{\|p_{1}-p_{j}\|}<\infty.$
Write $a_{j}=y_{j}+{\rm{i}}z_{j}=H(p_{j})$, and let $\{b_{1},...,b_{n},$...\} denote the set of distinct values in $\{a_{j}\}$. We assume that $\{b_{k}\}_{k=1}^{\infty}$ has no accumulation point on $\mathbb{C}$ and for all $k\geq 1$,
$m_{0}:=\#\{j\in\mathbb{N}^{+}:a_{j}=0\}<\infty,$
$m_{k}:=\#\{j\in\mathbb{N^{+}}:a_{j}=b_{k}\}<\infty.$
Relabel the monopoles as $\{p_{k,l}\}_{k,l}$ so that
$H(p_{k,l})\equiv b_{k},\ p_{k,l}=(x_{k,l},y_{k,l},z_{k,l}).$
Define an entire function
$P(u):=u^{m_{0}}\prod_{k:b_{k}\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}}\right)\right)^{m_{k}},$
where $c_{k}$ are properly chosen to ensure the infinite product to converge uniformly on every compact subset of $\mathbb{C}$. Let $S\subset\mathbb{C}^{3}$ be the hypersurface defined by
$u_{1}u_{2}=P(u_{3}).$
For $\{q_{j}\}_{j=1}^{\infty}=\{\pi^{-1}(p_{j})\}_{j=1}^{\infty}$, the charts that we take to cover $M_{0}:=M\setminus\{q_{j}\}_{j=1}^{\infty}$ are
$M^{-}:$ $=\pi^{-1}\left(\left\{u\neq b_{k}\text{ for all }k\right\}\cup\bigcup_{k}\left\{u=b_{k}\text{ and }x<x_{k,1}\right\}\right)$
$M^{+}:$ $=\pi^{-1}\left(\left\{u\neq b_{k}\text{ for all }k\right\}\cup\bigcup_{k}\left\{u=b_{k}\text{ and }x>x_{k,m_{k}}\right\}\right)$
$M_{k,l}:$ $=\pi^{-1}\left(\{u\neq b_{k}\text{ for all }k\}\cup\{u=b_{k}\text{ and }x_{k,l}<x<x_{k,l+1}\}\right)$
for $1\leq l\leq m_{k}-1$ and $k\geq 1$. Each chart is mapped by $H$ to an open subset of $\mathbb{C}$ and hence the base manifold is Stein. The $\mathbb{C}^{*}$-action makes each restriction $H:M_{\alpha}\to H(M_{\alpha})$ a trivial holomorphic principal $\mathbb{C}^{*}$-bundle.
We introduce coordinates $(u,v_{\alpha})$ on each $M_{\alpha}$ and describe the gluing in the form
$(u,v_{\alpha})\sim\left(u,f_{\alpha\beta}(u)v_{\beta}\right).$
The equivalence relation $f_{\alpha\beta}$ are expressed using integer-valued functions $J_{r}(\alpha)$:
$J_{r}(\alpha):=\left\{\begin{aligned} &0,\ &M_{\alpha}=M^{-},\\ &l,&M_{\alpha}=M_{r,l},\\ &m_{r},&M_{\alpha}=M^{+},\\ &0,&r\neq k\text{ on }M_{\alpha}=M_{k,l}.\end{aligned}\right$.
Then
$f_{\alpha\beta}(u)=u^{J_{0}(\alpha)-J_{0}(\beta)}\prod_{k:b_{k}\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}}\right)\right)^{J_{k}(\alpha)-J_{k}(\beta)}.$
Proof of Theorem 1.4.
For the surface $S\subset C^{3}$ defined by $u_{1}u_{2}=P(u)$, consider the map $\chi_{0}:M_{0}\to S\setminus\{0,0,b_{k}\}_{k}$:
$\chi_{0}|_{M^{-}}(u,v_{-})$ $=\left(\frac{P(u)}{v_{-}},v_{-},u\right),$
$\chi_{0}|_{M^{+}}(u,v_{+})$ $=\left(\frac{1}{v^{+}},P(u)v_{+},u\right),$
$\chi_{0}|_{M_{k,l}}(u,v_{k,l})$ $=\left(\frac{P(u)(u-b_{k})^{-l}}{v_{k,l}},(u-b_{k})^{l}v_{k,l},u\right).$
These three families of formulas agree on overlaps, so $\chi_{0}$ is a well-defined, holomorphic map. It extends holomorphically over the deleted points by setting $\chi(q_{k,l})=(0,0,b_{k})$, $q_{k,l}=\pi^{-1}(p_{k,l})$ and therefore defines a holomorphic map $\chi:M\to S$.
For each $k$ with $m_{k}\geq 2$ and $1\leq l\leq m_{k}-1$, consider the curve
$E_{k,l}:$ $=\overline{\{(u,v_{k,l})\in M_{k,l}:u=b_{k}\}},\ 1\leq l\leq m_{k}-1.$
By regarding $(u,v_{k,l})$ as $p_{k,l}$ when $v_{k,l}=0$ and as $p_{k,l+1}$ when $v_{k,l}=\infty$, we see that each $E_{k,l}$ is biholomorphic to $\mathbb{P}^{1}$ and their intersections satisfy
$E_{k_{1},l_{1}}\cap E_{k_{2},l_{2}}=\left\{\begin{aligned} &\{(u,v_{k,l_{1}})=(b_{k},\infty)\}=\{(u,v_{k,l_{2}})=(b_{k},0)\},&k_{1}=k_{2}=k,\ l_{2}=l_{1}+1,\\ &\emptyset,&k_{1}\neq k_{2}\text{ or }|l_{1}-l_{2}|\geq 2.\end{aligned}\right$.
For $k$ with $m_{k}\geq 2$, set
$E_{k}:$ $=\bigcup_{l=1}^{m_{k}-1}E_{k,l}=\chi^{-1}(0,0,b_{k})\subset M,$
$E:$ $=\bigcup_{k:m_{k}\geq 2}E_{k}.$
Then
$\chi^{-1}(0,0,b_{k})=E_{k},$
and the restriction $\chi|_{M\setminus E}:M\setminus E\to S\setminus\{(0,0,b_{k})\}_{k=1}^{n}$ is a biholomorphism. Since $\chi$ is proper, it follows that $\chi:M\to S$ is a resolution of singularities of $S$.
To compute the self-intersection of the exceptional curves, we apply the adjunction formula
$2g(E_{k,l})-2=\left(K_{M}+E_{k,l}\right)\cdot E_{k,l}.$
Here we have $g(E_{k,l})=0$ and $K_{M}\cdot E_{k,l}=0$ since $M$ is a hyperkähler manifold and $K_{M}$ is trivial. Hence $E_{k,l}^{2}=-2$ for all $k,l$. Thus every component of the exceptional divisor has self-intersection 2-2, and no component over any $b_{k}$ can be further contracted. Therefore $\chi:M\to S$ is a minimal resolution. ∎
Finally, recall that for a direction $v=(v_{1},v_{2},v_{3})\in S^{2}\subset\mathbb{R}^{3}$ we could choose complex structure $J_{v}=v_{1}J_{x}+v_{2}J_{y}+v_{3}J_{z}$ so that the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to $v$. If $v\in S^{2}$ is the direction satisfying that the projected points $a_{j}(v):=\Pi_{v}(p_{j})\in\mathbb{C}$ have pairwise distinct values $\{b_{k}(v)\}_{k}$ with no accumulation point on $\mathbb{C}$, and the multiplicity at each $b_{k}(v)$ is finite:
$m_{0}(v):=\#\{j\in\mathbb{N}^{+}:a_{j}(v)=0\}<\infty,$
$m_{k}(v):=\#\{j\in\mathbb{N^{+}}:a_{j}(v)=b_{k}(v)\}<\infty.$
Then we can define
$P_{v}(u):=u^{m_{0}(v)}\prod_{k:b_{k}(v)\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}(v)}\right)\right)^{m_{k}(v)}$
where $c_{k}(v)$ are properly chosen. For such a direction $v$, the above argument applies verbatim to $(M,J_{v})$, and show that $(M,J_{v})$ is biholomorphic to the minimal resolution of the hypersurface in $\mathbb{C}^{3}$ defined by $u_{1}u_{2}=P_{v}(u_{3})$, whose singularity type is $(A_{m_{1}-1},\cdots,A_{m_{k}-1},\cdots)$.